A Complete Bibliography of Publications in Scientometrics: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

activities [MGS79]. advance [Nar78]. allocation [BB79]. analysis [Gar79, KM79]. authorship [dBR79b].

bibliography [Vla78]. big [SM79]. Book [BH79, VP78]. bureaucracy [Sza79].

cancer [CS79]. carbides [BK79]. careers [KM79, Vla79c]. characteristics [MGS79]. chemistry [RLH79]. Citation [MM79, SM79, Gar79, RI79, SC79]. co [dBR79b]. co-authorship [dBR79b]. collaboration [dBR78, dBR79b, dBR79a]. computer [DRR79]. Cooperative [RLH79]. countries [RI79].

developed [RI79]. Differences [BB79]. disciplines [SC79]. distribution [Kun79]. distributions [Vla78]. domain [CS79]. Dynamics [BK79]. earth [BK79]. Editorial [de 78]. education [Vla79c]. employment [Vla79c]. evaluation [Gar79]. examination [SC79].

field [BK79]. flow [BK79]. Frequency [Vla78].

Gaps [IL79]. grants [BB79]. growth [CS79, Gil78].

history [dBR79b].

Identifying [HV78]. III [dBR79b]. indexes [SC79]. inequality [HV78]. information [BK79, Kun79, MGS79]. innovation [IL79]. input [IA78]. input-output [IA78]. international [DRR79]. inventories [IL79].

Knowledge [CS79].

Law [Vla78]. leading [MGS79]. legitimate [Gar79]. less [RI79]. little [SM79]. Lotka [Vla78].

matrix [KM79]. measurement [CS79]. measures [HV78]. Measuring [Gil78]. Mobility [Vla79a]. modern [dBR79b].

natural [dBR79b]. networks [DRR79]. News [Far78, Far79, FBV79, FB79a, FB79b]. Nobel [Vla79b].

Objectivity [Nar78]. options [DRR79]. orientation [BB79]. other [IL79]. output [IA78].

Part [dBR79b]. patent [Kun79]. patterns [MM79, SM79]. Penn [Vla79c]. performance [Vla78]. periphery [RI79]. petroleum [RLH79]. phenomena [Vla78]. Physics [Vla79c]. Prizes [Vla79b]. problem [CS79]. problems [Sza79]. Professionalization [dBR79b].

Quotations [Vla79d].

rare [BK79]. reflected [BB79]. related [Vla78]. relevance [Nar78]. Research [Sza79, BB79, BK79, DRR79, RLH79]. reviews [BH79, VP78]. revolutions [MM79].

Science [RI79, Gil78, HV78, IA78, KM79, SM79, SC79, Vla79a]. scientific [dBR78, dBR79b, CS79, KM79, MM79, Nar78, Vla78, dBR79a]. scientists [MGS79]. scientometrical [MGS79]. scientometrics [Vla79d]. set [HV78]. social [SC79]. some [Sza79]. special [KM79]. Specialties [SC79]. State [Vla79c]. statements [de 78]. structure [SC79]. structures [CS79]. Studies [dBR78, dBR79b, dBR79a, HV78, Nar78]. study [RI79]. system [IA78].

team [DRR79]. technology [IL79]. their [SC79]. three [RI79]. ties [KM79]. Time [Kun79]. tool [Gar79].

using [SC79].

versus [Nar78]. via [DRR79].

World [IA78].

References

Boalt:1979:DRO

[BB79] G. Boalt and U. Bergryd. Differences in research orientation reflected in the allocation of grants. Scientometrics, 1(2):151–159, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02016967.

Blaha:1979:BR

[BH79] K. Bláha and J. Hustopecký. Book reviews. Scientometrics, 1(2):195– 196, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016970.

Brozek:1979:DIF

[BK79] V. Brozek and P. Karen. Dynamics of information flow in the field of rare earth carbides research. Scientometrics, 1(4):339–357, May 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/BF02019305.

Chubin:1979:KSS

[CS79] D. E. Chubin and K. E. Studer. Knowledge and structures of scientific growth measurement of a cancer problem domain. Scientometrics, 1(2):171–193, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.com/

article/10.1007/BF02016969.

Beaver:1978:SSC

[dBR78] D. deB. Beaver and R. Rosen. Studies in scientific collaboration. Scientometrics, 1(1):65–84, September 1978. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016840.

deBeaver:1979:SSC

[dBR79a] D. B. de Beaver and R. Rosen. Studies in scientific collaboration. Scientometrics, 1(2):133–149, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016966.

Beaver:1979:SSC

[dBR79b] D. deB. Beaver and R. Rosen. Studies in scientific collaboration. Part III. Professionalization and the natural history of modern scientific co-authorship. Scientometrics, 1(3):231–245, March 1979. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016308.

deSollaPrice:1978:ES

[de 78] D. de Solla Price. Editorial statements. Scientometrics, 1(1):3–8, September 1978. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/10.
1007/BF02016836.

Dobrov:1979:NOT

[DRR79] G. M. Dobrov, R. H. Randolph, and W. D. Rauch. New options for team research via international computer networks. Scientometrics, 1(5–6):387–404, August 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.com/

article/10.1007/BF02016658.

Farkas:1978:N

[Far78] J. Farkas. News. Scientometrics, 1(1):101–105, September 1978. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016843.

Farkas:1979:Na

[Far79] J. Farkas. News. Scientometrics, 1(2):197–200, January 1979. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016971.

Farkas:1979:Nc

[FB79a] J. Farkas and J. Balázs. News. Scientometrics, 1(4):381–385, May 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02019308.

Farkas:1979:Nd

[FB79b] J. Farkas and J. Balázs. News. Scientometrics, 1(5–6):475–478, August 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02016663.

Farkas:1979:Nb

[FBV79] J. Farkas, Judith Balázs, and J. Vlachý. News. Scientometrics, 1(3): 285–293, March 1979. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016313.

Garfield:1979:CAL

[Gar79] E. Garfield. Is citation analysis a legitimate evaluation tool? Scientometrics, 1(4):359–375, May 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019306.

Gilbert:1978:MGS

[Gil78] G. Nigel Gilbert. Measuring the growth of science. Scientometrics, 1(1):9–34, September 1978. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.com/

article/10.1007/BF02016837.

Hustopecky:1978:ISI

[HV78] J. Hustopecký and J. Vlachý. Identifying a set of inequality measures for science studies. Scientometrics, 1(1):85–98, September 1978. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016841.

Inhaber:1978:WSI

[IA78] H. Inhaber and M. Alvo. World science as an input-output system. Scientometrics, 1(1):43–64, September 1978. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016839.

Inhaber:1979:GGT

[IL79] H. Inhaber and M. S. Lipsett. Gaps in “gaps in technology” and other innovation inventories. Scientometrics, 1(5–6):405–417, August 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/BF02016659.

Krauze:1979:MAS

[KM79] T. K. Krauze and R. McGinnis. A matrix analysis of scientific special ties and careers in science. Scientometrics, 1(5–6):419–444, August 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02016660.

Kunz:1979:TDP

[Kun79] M. Kunz. Time distribution of patent information. Scientometrics, 1(3):275–282, March 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.com/

article/10.1007/BF02016311.

Mulchenko:1979:SCI

[MGS79] Z. M. Mulchenko, Yu. V. Granovsky, and A. B. Strakhov. On scientometrical characteristics on information activities of leading scientists. Scientometrics, 1(4):307–325, May 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019303.

Moravcsik:1979:CPS

[MM79] M. J. Moravcsik and P. Murugesan. Citation patterns in scientific revolutions. Scientometrics, 1(2):161–169, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016968.

Narin:1978:OVR

[Nar78] F. Narin. Objectivity versus relevance in studies of scientific advance. Scientometrics, 1(1):35–41, September 1978. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02016838.

Rabkin:1979:SPC

[RI79] Y. M. Rabkin and H. Inhaber. Science on the periphery: a citation study of three less developed countries. Scientometrics, 1(3):261–274, March 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02016310.

Rabkin:1979:CRP

[RLH79] Y. M. Rabkin and J.-J. Lafitte-Houssat. Cooperative research in petroleum chemistry. Scientometrics, 1(4):327–338, May 1979. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02019304.

Small:1979:SDS

[SC79] H. G. Small and D. Crane. Specialties and disciplines in science and social science: An examination of their structure using citation indexes. Scientometrics, 1(5–6):445–461, August 1979. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02016661.

Shearer:1979:CPL

[SM79] E. Shearer and M. J. Moravscik. Citation patterns in little science and big science. Scientometrics, 1(5–6):463–474, August 1979. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016662.

Szalai:1979:RRS

[Sza79] A. Szalai. Research on research and some problems of research bureaucracy. Scientometrics, 1(3):247–260, March 1979. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02016309.

Vlachy:1978:FDS

[Vla78] J. Vlachý. Frequency distributions of scientific performance a bibliography of Lotka’s Law and related phenomena. Scientometrics, 1 (1):107–130, September 1978. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.com/

article/10.1007/BF02016844.

Vlachy:1979:MS

[Vla79a] J. Vlachý. Mobility in science. Scientometrics, 1(2):201–228, January 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.1007/
BF02016972.

Vlachy:1979:NP

[Vla79b] J. Vlachý. Nobel Prizes. Scientometrics, 1(3):295–301, March 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/BF02016314.

Vlachy:1979:PCE

[Vla79c] J. Vlachý. Physics careers, employment and education (penn state, 1977). Scientometrics, 1(3):283–284, March 1979. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/BF02016312.

Vlachy:1979:QS

[Vla79d] J. Vlachý. Quotations and scientometrics. Scientometrics, 1(4):377– 380, May 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/10.
1007/BF02019307.

Vlachy:1978:BR

[VP78] J. Vlachý and V. Polácek. Book reviews. Scientometrics, 1(1):99–100, September 1978. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016842.